OpenStudy (anonymous):

Which factor do you cancel from the numerator and denominator to simplify (x2-3x-4)/(x2-16) A x + 3 B x − 4 C x − 3 D x + 4 I think its D, is that right?

OpenStudy (anonymous):

D is correct

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

if i were you i would have done B :P

OpenStudy (anonymous):

can I ask why?

OpenStudy (anonymous):

yes you can. we need to simplify first one the factors are factors are (x-4)(x+1) now for second one it is x2-b2=(x-b)(x+b) x2-16= (x-4)(x+4) so x-4 is present in numerator and denominator so thats the correct answer.

OpenStudy (anonymous):

ohhh, okay. I see!

OpenStudy (anonymous):

Im gonna ask one more, but ill post it here instead of doing in individually. Enrique has two rectangular dive flags. The flags are similar and have dimensions, in inches, as shown in the figure below. one rectangle 3x by 4x-8 and a smaller rectangle 16 by 15 What is the value, in inches, of x? A 6 B 9 C 10 D 12

OpenStudy (anonymous):

can anyone help with the second one?

OpenStudy (anonymous):

x=10

OpenStudy (anonymous):

3x/15=4x-8/16 48x=60x-120 x=10 :D

OpenStudy (anonymous):

ohhh, okay thanks again! :)

OpenStudy (anonymous):

at least i deserve a medal :P

OpenStudy (anonymous):

I already gave you one, I cant give another or I would :)